let a and b be sets. show that (a∩b)⊆a


1
Expert's answer
2022-03-14T08:15:16-0400

Solution


Consider the two sets are


"a = \left\{ {1,2,3,4,5,6,7} \right\}\" and "b = \left\{ {5,6,7,8,9} \right\}\"


Then


"\left( {a \cap b} \right) = \left\{ {1,2,3,4,5,6,7} \right\} \cap \left\{ {5,6,7,8,9} \right\}\\"


"\left( {a \cap b} \right)
 = \left\{ {5,6,7} \right\}"


We can see all the elements "5, 6, 7" of set the "\left( {a \cap b} \right)" are also present in the set "a = \left\{ {1,2,3,4,5,6,7} \right\}\"


Hence the set "a \cap b\"is a subset of the set "a = \left\{ {1,2,3,4,5,6,7} \right\}\"


We can write it mathematically as


"\left( {a \cap b} \right) \subseteq a\"